# Multi-messenger lensing time delay as a probe of the graviton mass

### v1

@@ VI Conclusion
In this work, we have investigated the effects of a massive graviton on a lensed gravitational wave, and the potential for a lensed multi-messenger event to constrain the mass of the graviton. After illustrating how a massive graviton affects the dispersion relation, we investigated three aspects that change in the presence of a massive graviton, namely, geodesics, the time delays between different images, and the magnification of the signals due to a gravitational lens. We computed the first order corrections to these quantities in powers of $m^{2}/\omega^{2}$.
@@
Starting from the dispersion relation, we derived an expression for the time delay between different lensed signals for a massive graviton, which differs from the massless case by a factor of $1-m^{2}/(2\omega^{2})$. Accounting for the different scattering angles of massive gravitons with respect to massless photons led to two extra terms which turn out to cancel. This is because while the massive photons travel a little bit longer geometrically than photons, because of their larger scattering angle, this is compensated by experiencing less time dilation from the lens. This is due to the impact parameter being larger than for massless photons, as we illustrate in Fig. 2. This cancellation makes the difference in time delay between photons and massive gravitons arise solely because of the different group velocity of the gravitons with respect to massless photons, as may be understood from Eq. (46). We argued that we could constrain the mass of the graviton by comparing GW and EM time delays between different images. This effectively cancels the contributions that depend on cosmology and the lens model, as should be clear from Eq. 46, such that a fully model-independent constraint can be imposed on the mass of the graviton. We find that for a merger in the mHz which is relevant for LISA, the constrain can reach $m<3\cdot 10^{-23}$eV/c2. Note that, additionally, the constraint is independent of the waveform model, which is not the case for the dispersion constraints.
@@
Finally, microlensing by stars in the lens could in principle affect the EM and GW signals causing a discrepancy in both magnification and time delays. This holds provided that the geometric optics approximation remains valid, which is less obvious for GWs than it is for EM waves, given the wavelenghts involved. In such a scenario, since the geodesics followed by massive gravitons and light are slightly different, there is a possibility that one of the two gets micro-magnified but not the other. Such a signal could be used as a smoking gun signature of a non-zero graviton mass and we leave this possibility to future work. As for the time delay, microlensing of the EM signal effectively adds an extra delay term to the left hand side of Eq.(46). This is however of order $10^{-5}$s [59, 60], much smaller than both $\Delta t_{\gamma}$ and $\sigma_{\Delta t_{\gamma}}$, meaning that our estimate on the graviton mass constraints would remain unaffected.
Overall, we have shown that a single multi-messenger lensed event can probe the graviton mass in a fully model-independent fashion and constrain $m<3\cdot 10^{-23}$ eV/c2 using the time delay between different images.
We have mentioned that lensed gravitational wave signals are expected to be rare, with only one observed multiply-imaged signal every $\sim 1500$ detections. While the current number of GW detections and the absence of comprehensive all-sky coverage in the EM domain make the detection of a golden event currently unlikely, the methods described in this paper may be used in the foreseeable future. Over the coming decade, gravitational-wave astronomy will see significant advancements with the advent of LISA, Einstein Telescope, and Cosmic Explorer. The latter two in particular forecast detections in the thousands [61, 62], and with new surveys such as LSST monitoring large portions of the sky, there is a chance to observe a multi-messenger strongly lensed event, which will enable a fully model-independent probe of the graviton mass.

### v2

@@ VI Conclusion
In this work, we have investigated the effects of a massive graviton on a lensed gravitational wave, and the potential for a lensed multimessenger event to constrain the mass of the graviton. After illustrating how a massive graviton affects the dispersion relation, we investigated three aspects that change in the presence of a massive graviton, namely, geodesics, the time delays between different images, and the magnification of the signals due to a gravitational lens. We computed the first-order corrections to these quantities in powers of $m^{2}/\omega^{2}$.
@@
Starting from the dispersion relation, we derived an expression for the time delay between different lensed signals for a massive graviton, which differs from the massless case by a factor of $1-m^{2}/(2\omega^{2})$. Accounting for the different scattering angles of massive gravitons with respect to massless photons led to two extra terms which turn out to cancel. This is because while the massive gravitons travel a little bit longer geometrically than photons, because of their larger scattering angle, this is compensated by experiencing less time dilation from the lens. This is due to the impact parameter being larger than for massless photons, as we illustrate in Fig. 2. This cancellation makes the difference in time delay between photons and massive gravitons arise solely because of the different group velocity of the gravitons with respect to massless photons, as may be understood from Eq. (42). We argued that we could constrain the mass of the graviton by comparing GW and EM time delays between different images. This effectively cancels the contributions that depend on cosmology and the lens model, as should be clear from Eq. (42), such that a fully model-independent constraint can be imposed on the mass of the graviton. We find that for a merger in the milli-Hertz which is relevant for LISA, the constraint can reach $m<3\cdot 10^{-23}$eV/c2 when lensed by a cluster-scale lens. Note that, additionally, the constraint is independent of the waveform model, which is not the case for the dispersion constraints.
@@
Finally, microlensing by stars in the lens could in principle affect the EM and GW signals causing a discrepancy in both magnification and time delays. This holds provided that the geometric optics approximation remains valid, which is less obvious for GWs than it is for EM waves, given the wavelengths involved. In such a scenario, since the geodesics followed by massive gravitons and light are slightly different, there is a possibility that one of the two gets micromagnified but not the other. Such a signal could be used as a smoking gun signature of a nonzero graviton mass, and we leave this possibility to future work. As for the time delay, microlensing of the EM signal effectively adds an extra delay term to the right-hand side of Eq. (40). This is, however, of order $10^{-5}$ s [48, 53], much smaller than both $\Delta t_{\gamma}$ and $\sigma_{\Delta t_{\gamma}}$, meaning that our estimate on the graviton mass constraints would remain unaffected. Note that a graviton charge would be partially degenerate with a graviton mass, as discussed in [56].
Overall, we have shown that a single multimessenger lensed event can probe the graviton mass in a fully model-independent fashion and constrain $m<3\cdot 10^{-23}$eV/c2 using the time delay between different images.
We have mentioned that lensed gravitational wave signals are expected to be rare, with only one observed multiply imaged signal every $\sim 1500$ detections. While the current number of GW detections and the absence of comprehensive all-sky coverage in the EM domain make the detection of a golden event currently unlikely, the methods described in this paper may be used in the foreseeable future. Over the coming decade, gravitational-wave astronomy will see significant advancements with the advent of LISA, Einstein Telescope, and Cosmic Explorer. The latter two in particular forecast detections in the thousands [51, 33], and with new surveys such as LSST monitoring large portions of the sky, there is a chance to observe a multimessenger strongly lensed event, which will enable a fully model-independent probe of the graviton mass.
@@ VII Acknowledgements
VIII Data Availability
No data were created or analyzed in this study.